# Variational simulation of the Lipkin-Meshkov-Glick model on a neutral atom quantum computer

### v2

@@ I Introduction
Scaling up of digital quantum processors for running large circuits requires establishing connectivity between distant qubits. Universal paradigms for enabling arbitrary connectivity rely on either moving quantum states across the processor with a sequence of swap gates[1] or using teleportation[2]. Both of these approaches incur additional resource overhead as long swap chains increase circuit depth, and efficient teleportation depends on the availability of high-fidelity resource states. Physical motion of qubits is a third alternative that was originally introduced for the trapped ion platform[3, 4, 5], and has been successfully implemented with neutral atom qubits for logical processing[6, 7, 8], as well as in atomic tweezer clocks[9]. In this work, we incorporate both atom transport for dynamical connectivity and rapid optical beam scanning for site-specific optical control for simulation of the Lipkin-Meshkov-Glick(LMG) model using the Variational-Quantum-Eigensolver(VQE) algorithm. We show that application of efficient encoding strategies together with noise mitigation techniques[10] enables finding ground-state energies with an accuracy beyond that supported by the intrinsic physical gate fidelities.
@@ IV.1 Three spin encodings
To improve the ground-state energy estimation, we switched to a different qubit encoding scheme based on a Gray code encoding as described in Appendix A. This encoding technique uses qubits to encode spins much more efficiently, allowing us to encode three spins using a single qubit (see Appendix A.2). The corresponding three-spin VQE circuit is composed of single-qubit gates with the same variational parameter $\theta$ that was used in the individual spin encoding scheme. This simplification allows improved ground-state energy estimation. In the Gray code basis, the ground-state energy can be determined by making measurements in the $\sf X$ and $\sf Z$ bases as explained in Appendix A.2 using the weights in Table A1. We determined the energy as a function of $\theta$ and fit to a cosine function (see Fig. 3b) to find a ground-state energy of -2.43(5) (much closer to the theoretical value of -2.5). A comparison of these two encoding techniques is summarized in Table 1.
@@ IV.2 More than three spins
Three qubits were needed to encode the ansatz states for 9 and 15 spins. All three-qubit VQE algorithms use the same general quantum circuit with three $\sf CZ$ gates (see Fig. 6). Both spin numbers also require the same Pauli string measurements ($\sf ZZZ$, $\sf XZZ$, $\sf ZXZ$, and $\sf ZZX$) to determine the energy expectation of the LMG Hamiltonian using the weights in Tables A4 and A5. The ansatz states for 9 and 15 spins required 4 and 7 variational parameters, respectively, so we used the Nelder-Mead algorithm to find the optimal variational parameter settings, as we did for the 7-spin VQE optimization. The optimization protocol did not fully converge to the optimal variational parameter settings for either spin number. The 9-spin optimization converged to a variational parameter set which corresponded to a theoretical energy expectation 23% higher than the theoretical ground-state energy (see Fig. 7a-e). We further refined the 9-spin variational parameter settings with 1D line scans of each of the four angles and fit the energies of each set to a parabola (see Fig. 8). The set of variational parameters from these fits corresponded to a theoretical energy value within 3.3% of the ground-state energy. The larger parameter number and more complicated optimization landscape prevented the 15-spin VQE algorithm from converging to a variational parameter set near the theoretical optimum (the theoretical value of the converged parameter set was 48% higher than the theoretical ground state energy); this large error prevented us from finding an accurate ground-state energy estimate for 15 spins. Further discussion of the 15-spin VQE optimization is found in Appendix C.
@@ V Summary
The agreement of the simulated ground-state energies with exact LMG model results was improved by using ZNE for error mitigation. Error mitigation with ZNE provided close agreement between the VQE results and theory for up to 7 spins and agreement to about 5% for 9 spins. Attempts to extend to a 15-spin problem with 7 variational parameters resulted in poor convergence. We attribute the lack of convergence to a combination of noisy gate execution and the possible existence of barren plateaus in the high-dimensional state space. One solution will be full quantum error correction, which requires many more qubits and higher $\sf CZ$ fidelity than was available here. Alternatively, it will be of interest to consider an intermediate approach that goes beyond error mitigation with small error detecting codes. Such codes can provide improved results with only a moderate increase in qubit overhead, as has recently been demonstrated for quantum chemistry and quantum materials simulations[30, 31].
@@ A.2 Gray code basis
The general formula for arbitrary $N$ is
@@
$+\prod_{l=0}^{\lfloor J\rfloor}\sin(\theta_{l+1})\ket{g_{\lfloor J\rfloor}}.$ (17)
@@ Appendix C Simulation of 15 spin Hamiltonian
We attempted to simulate the ground state of $N=15$ spins encoded in 3 qubits with Gray code encoding. In order to simulate 15 spins, seven variational parameters are required, and the necessary basis transformations are the four Pauli strings bases: $\sf ZZZ,XZZ,ZXZ,ZZX$. The results are shown in Fig. 9. We observed that the variational angles start to converge, but except for $\theta_{3}$, all variational angles (a-g) deviated from the theoretically calculated angles(shown with dashed lines). The energy in panel (h) also showed very weak convergence towards the ground state.
The lack of convergence for $N=15$ can be primarily attributed to the dimensionality of the search space. The $N=15$ circuit shown in Fig. 6 has the same width and gate depth as for the $N=9$ case but requires searching in a seven-dimensional space as opposed to the four-dimensional space for $N=9$. The convergence for $N=9$ was poor and required linear parameter scans and ZNE as seen in Figs. 7 and 8 and Table 1. The $N=15$ performance was significantly worse, so we did not attempt ZNE. The degradation of convergence for the same circuit complexity may be due to so-called barren plateaus that can be pervasive when the problem dimensionality increases in the presence of noise[35].
@@ Appendix E Experimental Apparatus
During circuit execution, the atoms were reconfigured mid-circuit by ramping the frequencies with which the 1064-nm array’s AODs were driven (see Appendix F). This reconfiguration allowed gate operation provided by a combination of: 1. Microwaves for global $\sf{R}_{\phi}$ gates (where the rotation axis is located in the $x-y$ plane of the Bloch sphere at an angle $\phi$ from the $x$-axis). 2. a 459-nm beam (Vexlum VALO-SHG-SF) blue-detuned from the $7p_{1/2}$ level by 1.05 GHz to provide local $\sf Z$-rotations through the AC Stark shift. 3. the same 459-nm beam plus a 1040-nm beam (M-Squared SolsTiS-SA-PSX-XL) to provide $\sf CZ$-gates. Together, these lasers could drive Rabi oscillations at up to $2\pi\times 6.5$ MHz; however, a lower Rabi frequency of $2\pi\times 1.7$ MHz was found to give better gate fidelity. We hypothesize that off-resonant Raman transitions out of the qubit basis were responsible for the lower fidelity at high Rabi frequencies. More details about the universal gate set can be found in the main text.
@@
Figure 11: Comparison of Gaussian and super Gaussian beams with $p=2.6$ and waists $w=8.6~{}\mu$m (see 18). a) and b) show the normalized intensity and its derivative of a Gaussian vs super Gaussian as a function of position; the super Gaussian beam has a more uniform intensity distribution at $1.25~{}\mu$m from the beam center where the atoms are trapped. c) Simulation of a Stark shifted Ramsey experiment showing reduced dephasing with a super-Gaussian beam. Parameters: 1064 nm trap with waist $w=1.2~{}\mu\rm m$, depth $k_{B}\times 600~{}\mu\rm K$, and atom temperature $15~{}\mu$K. The atom position was displaced $d=1.25~{}\mu$m from the beam centers.
@@ Appendix F Mid-circuit atom reconfiguration
For a wide range of parameters, the minimal jerk profile provides less motional heating for the same transport time. Using the parameters given above, we find $\delta n_{\rm cj}=0.053$ and $\delta n_{\rm mj}=5.7\times 10^{-4}$, almost two orders of magnitude smaller. With these parameters, less than $1\mu$K of heating was observed after 5 round-trip ramps. The qubit coherence was preserved by applying Hahn-echo pulses between round-trip transports to cancel accumulated phase shifts.
Figure 12: Quantum gates fidelity characterizations using Randomized benchmarking (RB). a) Shows the RB circuit used to measure $\sf CZ$ gate fidelity. b) Shows the exponential decay of the two-atom ground state population as increasing number of $\sf CZ$ gates ($\sf CZ_{(q0,q1)}$ on the left and $\sf CZ_{(q1,q2)}$ on the right). c) Shows the standard Clifford single qubit RB circuit used to characterize the global and local single-qubit gate sets. d) Shows the results of global microwave RB of all three sites. e) Shows the results of RB of the local single qubit gates on all three sites.
@@ Appendix G Randomized benchmarking of quantum gates set
We used randomized benchmarking (RB) to characterize the performance of the quantum gate set. For the $\sf CZ$ gate fidelity characterization, we use the symmetric interleaved RB sequence demonstrated in ref. [24]. Up to 12 $\sf CZ$ gates were interleaved with random single-qubit rotations applied globally with microwaves (see Fig. 12a). The $\sf CZ$ gates are performed in pairs with an $\sf X$ gate inserted between to eliminate the single-qubit frame shift acquired during each $\sf CZ$ gate. We fit the data with an exponential decaying function to extract a gate fidelity of $\mathcal{F}_{{\sf CZ}(0,1)}=97.7(1)\%$ for qubit pair $q_{0}$ and $q_{1}$ and $\mathcal{F}_{{\sf CZ}(1,2)}=96.4(2)\%$ for qubit pair $q_{1}$ and $q_{2}$.
We used standard single-qubit Clifford randomized benchmarking [39] to characterize both global and local single-qubit gates (see Fig. 12c for a typical circuit). For global single-qubit rotations, we apply a global variant of the standard randomized benchmarking as shown in [40] (see Fig.12d). After fitting, we extracted an average gate fidelity of 0.99989(1). The faster decay after 500 quantum gates is due to changes in the qubit frequency due to a transient increase of the magnetic field. A similar RB procedure was used to find the gate fidelity of a set of local Clifford gates constructed using global $\sf{R}_{\phi}$ and local $\sf{R}_{z}$ gates (see Fig.12e). We measured an average local Clifford gate fidelity of 0.99941(4).
@@ Appendix H Optimized $\sf CZ$ gate
In this section, we present the experimental modifications we made to improve the $\sf CZ$ gate fidelity. These modifications were made after the VQE measurements were performed and highlight some strategies for fidelity improvement. These modifications include: 1. We excited to a lower Rydberg level, $55s_{1/2}$, which has a larger dipole matrix element compared to $66s_{1/2}$. The larger matrix element reduces the intensity required to drive the Rydberg transition at a given Rydberg Rabi rate. This effectively reduces infidelity caused by AC Stark shift fluctuations resulting from intensity fluctuations. 2. We found the optimal Rydberg Rabi frequency to be $2\pi\times 1$ MHz (versus $2\pi\times 1.7$ MHz in the VQE experiment). Although gates performed at lower Rabi frequency are more susceptible to some error sources such as Rydberg decay, the increase in such error sources is more than compensated for by the reduction in dephasing during the Rydberg excitation that was observed in the experiment. This dephasing reduction is not fully understood, but we believe that it is related to intensity noise found in the $459$-nm Rydberg laser. Additional measurements and simulations are needed to confirm this hypothesis. 3. After optical pumping, we added a trap power ramp to reduce the depth from $600\mu$K to $260~{}\mu$K. This ramping, along with improved cooling parameters, reduced the atom temperature to $6~{}\mu$K. After these changes, we measured a $\sf CZ$ gate fidelity of $\mathcal{F}_{\sf CZ}=0.986(1)$ (see Fig.13) using the methods described in Appendix G.

### v3

@@ I Introduction
Scaling up of digital quantum processors for running large circuits requires establishing connectivity between distant qubits. Universal paradigms for enabling arbitrary connectivity rely on either moving quantum states across the processor with a sequence of swap gates[1] or using teleportation[2]. Both of these approaches incur additional resource overhead as long swap chains increase circuit depth, and efficient teleportation depends on the availability of high-fidelity resource states. Physical motion of qubits is a third alternative that was originally introduced for the trapped ion platform[3, 4, 5], and has been successfully implemented with neutral atom qubits for logical processing[6, 7, 8], as well as in atomic tweezer clocks[9]. In this work, we demonstrate several innovations for quantum circuit operation with neutral atom qubits. Using both atom transport for dynamical modification of the connectivity map and rapid optical beam scanning for site-specific optical control we simulate the Lipkin-Meshkov-Glick(LMG) model using the Variational-Quantum-Eigensolver(VQE) algorithm. We show that application of efficient encoding strategies together with noise mitigation techniques[10] enables finding ground-state energies with an accuracy beyond that supported by the intrinsic physical gate fidelities.
@@ IV.1 Three spin encodings
To improve the ground-state energy estimation, we switched to a different qubit encoding scheme based on a Gray code encoding as described in Appendix A. This encoding technique uses qubits to encode spins much more efficiently, allowing us to encode three spins using a single qubit. As is explained in detail in Appendix A.2 the ansatz state for $N$ spins has $J=N/2$ variational parameters $\theta_{j}$ and takes the form
$+\prod_{l=0}^{\lfloor J\rfloor-1}\sin(\theta_{l+1})\ket{g_{\lfloor J\rfloor}}.$ (2)
where $\Theta=\theta_{1},\theta_{2},...\theta_{J}$.
The corresponding three-spin VQE circuit is composed of single-qubit gates with the same variational parameter $\theta$ that was used in the individual spin encoding scheme. This simplification allows improved ground-state energy estimation. In the Gray code basis, the ground-state energy can be determined by making measurements in the $\sf X$ and $\sf Z$ bases as explained in Appendix A.2 using the weights in Table A1. We determined the energy as a function of $\theta$ and fit to a cosine function (see Fig. 3b) to find a ground-state energy of -2.43(5) (much closer to the theoretical value of -2.5). A comparison of these two encoding techniques is summarized in Table 1.
@@ IV.2 More than three spins
Three qubits were needed to encode the ansatz states for 9 and 15 spins. All three-qubit VQE algorithms use the same general quantum circuit with three $\sf CZ$ gates (see Fig. 6). Both spin numbers also require the same Pauli string measurements ($\sf ZZZ$, $\sf XZZ$, $\sf ZXZ$, and $\sf ZZX$) to determine the energy expectation of the LMG Hamiltonian using the weights in Tables A4 and A5. The ansatz states for 9 and 15 spins required 4 and 7 variational parameters, respectively, so we used the Nelder-Mead algorithm to find the optimal variational parameter settings, as we did for the 7-spin VQE optimization. The optimization protocol did not fully converge to the optimal variational parameter settings for either spin number. The 9-spin optimization converged to a variational parameter set which corresponded to a theoretical energy expectation 23% higher than the theoretical ground-state energy (see Fig. 7a-e). We further refined the 9-spin variational parameter settings with 1D line scans of each of the four angles and fit the energies of each set to a parabola (see Fig. 8). The set of variational parameters from these fits corresponded to a theoretical energy value within 3.3% of the ground-state energy. The larger number of parameters and more complicated optimization landscape prevented the 15-spin VQE algorithm from converging to a variational parameter set near the theoretical optimum (the theoretical value of the converged parameter set was 48% higher than the theoretical ground state energy); this large error prevented us from finding an accurate ground-state energy estimate for 15 spins. Further discussion of the 15-spin VQE optimization is found in Appendix C.
@@ V Summary
The agreement of the simulated ground-state energies with exact LMG model results was improved by using ZNE for error mitigation. Error mitigation with ZNE provided close agreement between the VQE results and theory for up to 7 spins and agreement to about 5% for 9 spins. Attempts to extend to a 15-spin problem with 7 variational parameters resulted in poor convergence. As is shown in Appendix C the lack of convergence is due not just to gate execution errors, but also results from the possible existence of barren plateaus in the high-dimensional state space and sensitivity to shot-noise from the limited number of bitstring measurements allocated to each circuit. These factors resulted in the Nelder-Mead optimizer getting trapped in local minima without finding the global optimal solution.
Possible solutions may involve an expanded search over multiple initial conditions together with dynamic allocation of the number of iterations used for each circuit. To reduce the impact of gate errors full quantum error correction is ultimately the goal, but this will require many more qubits and higher $\sf CZ$ fidelity than was available here. Alternatively, it will be of interest to consider an intermediate approach that goes beyond error mitigation with small error detecting codes. Such codes can provide improved results with only a moderate increase in qubit overhead, as has recently been demonstrated for quantum chemistry and quantum materials simulations[30, 31].
@@ A.2 Gray code basis
The general formula for arbitrary $N{\color[rgb]{0,0,0}\geq 2}$ is
@@
$+\prod_{l=0}^{\lfloor J\rfloor{\color[rgb]{0,0,0}-1}}\sin(\theta_{l+1})\ket{g_{\lfloor J\rfloor}}$ (18)
with $J=N/2$ and $\Theta=\theta_{1},\theta_{2},...\theta_{J}$.
@@ Appendix C Simulation of 15 spin Hamiltonian
We attempted to simulate the ground state of $N=15$ spins encoded in 3 qubits with Gray code encoding. In order to simulate 15 spins, seven variational parameters are required, and the necessary basis transformations are the four Pauli string bases: $\sf ZZZ,XZZ,ZXZ,ZZX$. The results are shown in Fig. 9. We observed that the variational angles begin to converge, but except for $\theta_{3}$, all variational angles (a-g) deviated from the theoretically calculated angles(shown with dashed lines). The energy in panel (h) also showed very weak convergence towards the ground state.
The lack of convergence for the $N=15$ case is due to a number of factors including the dimensionality of the search space. The $N=15$ circuit shown in Fig. 6 has the same width and gate depth as for the $N=9$ case but requires searching in a seven-dimensional space as opposed to the four-dimensional space for $N=9$. The convergence for $N=9$ was poor and required linear parameter scans and ZNE as seen in Figs. 7 and 8 and Table 1. As demonstrated by the improvement in performance of the $N=9$ qubit case by ZNE, noise is another contributing factor. In the context of the VQE algorithm, there are two sources of noise. One is a consequence of imperfect quantum circuits and measurements. The second is due to the shot-noise inherent to the calculation of the expectation value for a given set of variational parameters. This shot-noise scales with the inverse square root of the number of bitstring measurements allocated to each circuit. We see a clear indication that circuit noise can cause a systematic difference between measured energy values and true values in the drift between the experimental and theory curves in later optimization iterations in Fig. 7, which we expect is due to gate calibration drift. Gate calibration drift imposes constraints on the maximum iteration depth we can implement, as well as the number of bitstring measurements allocated to each set of variational parameters.
Figure 10 shows the result of optimizing the $N=15$ case in the absence of both circuit noise and shot noise by directly evaluating the expectation value classically. Even in the absence of noise, the optimization did not converge to the optimal value of -53.47. The optimization was terminated at an iteration depth of 100, in excess of the iteration depths that each of the quantum optimizations were terminated due to run-time constraints imposed by gate calibration drift. To illustrate the consequence of shot noise at the level of 400 measurements per circuit used in the quantum simulations, we include error bars showing a one standard deviation range about the true mean. Even at one standard deviation, the shot noise would be large relative to the difference in energy values of adjacent iterations, which dramatically limits an optimizer’s ability to make progress.
In this noise-free optimization, each variational parameter was initialized to 0 to match the initial conditions of the quantum optimization. With an allowance for a greater maximum iteration depth, Nelder-Mead makes additional progress, but still fails to converge to the global optimum. This suggests that the outcome is sensitive to the initial variational parameters and the choice of the classical optimization algorithm. Collectively, these considerations imply that improved performance would involve some combination of dynamically allocating bitstring measurements until enough strings have been gathered to reduce the standard error or the measurement below a desired threshold, running multiple optimizations with randomized initial variational parameters, and a potentially different choice of classical algorithm. These improvements come with increases in real-world runtime of the algorithm and their benefits must be considered accordingly.
Figure 10: Results for $N=15$ spins by direct classical evaluation of the energy of the ansatz state during Nelder-Mead optimization. The optimizer does not find the true ground state (dashed line) within the allocated number of iterations, even in the absence of circuit noise or measurement shot noise.
@@ Appendix E Experimental Apparatus
During circuit execution, the atoms were reconfigured mid-circuit by ramping the frequencies with which the 1064-nm array’s AODs were driven (see Appendix F). This reconfiguration allowed gate operation provided by a combination of: 1. Microwaves for global $\sf{R}_{\phi}$ gates (where the rotation axis is located in the $x-y$ plane of the Bloch sphere at an angle $\phi$ from the $x$-axis). 2. a 459-nm beam (Vexlum VALO-SHG-SF) blue-detuned from the $7p_{1/2}$ level by 1.05 GHz to provide local $\sf Z$-rotations through the AC Stark shift.  [^1] Note that the 1064-nm traps are left on during local $\sf Z$-gates. The traps shift both $6s_{1/2}$ and $7p_{1/2}$. However, the magnitude of this shift is small compared to the detuning of the 459-nm light from this transition, so the only way that such shifts affect the gates is to give a small change in gate duration. This affect is included in the gate calibration. 3. The same 459-nm beam and a 1040-nm beam (M-Squared SolsTiS-SA-PSX-XL) provide $\sf CZ$-gates. Together, these lasers could drive Rabi oscillations at up to $2\pi\times 6.5$ MHz; however, a lower Rabi frequency of $2\pi\times 1.7$ MHz was found to give better gate fidelity. We hypothesize that off-resonant Raman transitions out of the qubit basis were responsible for the lower fidelity at high Rabi frequencies. More details about the universal gate set can be found in the main text.
@@
Figure 12: Comparison of Gaussian and super Gaussian beams with $p=2.6$ and waists $w=8.6~{}\mu$m (see 19). a) and b) show the normalized intensity and its derivative of a Gaussian vs super Gaussian as a function of position; the super Gaussian beam has a more uniform intensity distribution at $1.25~{}\mu$m from the beam center where the atoms are trapped. c) Simulation of a Stark shifted Ramsey experiment showing reduced dephasing with a super-Gaussian beam. Parameters: 1064 nm trap with waist $w=1.2~{}\mu\rm m$, depth $k_{B}\times 600~{}\mu\rm K$, and atom temperature $15~{}\mu$K. The atom position was displaced $d=1.25~{}\mu$m from the beam centers.
@@ Appendix F Mid-circuit atom reconfiguration
For a wide range of parameters, the minimal jerk profile provides less motional heating for the same transport time. Using the parameters given above, we find $\delta n_{\rm cj}=0.053$ and $\delta n_{\rm mj}=5.7\times 10^{-4}$, almost two orders of magnitude smaller. With these parameters, less than $1\mu$K of heating was observed after 5 round-trip ramps. The qubit coherence was preserved by applying two Hahn-echo pulses between round-trip transports to cancel accumulated phase shifts. The first Hahn echo occurs after the Rydberg pulse, while the second takes place after the second transport. Each Hahn echo is a global microwave X rotation, with each pulse lasting 8 $\mu$s. The entire sequence spans approximately 650 $\mu$s. In addition to canceling the accumulated phase shifts, the Hahn echo pulses mitigate the qubit dephasing during the transport sequence.
Figure 13: Quantum gate fidelity characterizations using Randomized benchmarking (RB). a) Shows the RB circuit used to measure $\sf CZ$ gate fidelity. b) Shows the exponential decay of the two-atom ground state population as increasing number of $\sf CZ$ gates ($\sf CZ_{(q0,q1)}$ on the left and $\sf CZ_{(q1,q2)}$ on the right). The population at 0 $\sf CZ$ gates is limited primarily by SPAM and errors on the target two-atom state due to imperfect microwave rotations. c) Shows the standard Clifford single qubit RB circuit used to characterize the global and local single-qubit gate sets. d) Shows the results of global microwave RB of all three sites. $r_{c}=(1-\mathcal{F}_{c})$ is the average error per Clifford gate, extracted from the decay fit [39]. e) Shows the results of RB of the local single qubit gates on all three sites.
@@ Appendix G Randomized benchmarking of quantum gates set
We used randomized benchmarking (RB) to characterize the performance of the quantum gate set. After applying a random gates sequence, followed by a final correction gate, the population in the |$\ket{0}$ state $P(\ket{0})_{m}$, is fit to an exponential decaying function $P(\ket{0})_{m}=A_{0}p^{m}+B_{0}$ where $m$ is the number of gates. The averaged error per Clifford gate $r_{c}$, is given by [39]
$r_{c}=1-\mathcal{F}_{c}=\frac{(d-1)(1-p)}{d}$ (20)
where $d=2^{n}$ is the dimension of the $n$ qubit system.
For the $\sf CZ$ gate fidelity characterization, we use the symmetric interleaved RB sequence demonstrated in ref. [24]. Up to 12 $\sf CZ$ gates were interleaved with random single-qubit rotations applied globally with microwaves (see Fig. 13a). The $\sf CZ$ gates are performed in pairs with an $\sf X$ gate inserted between to eliminate the single-qubit frame shift acquired during each $\sf CZ$ gate. We fit the data with an exponential decaying function to extract a gate fidelity of $\mathcal{F}_{{\sf CZ}(0,1)}=97.7(1)\%$ for qubit pair $q_{0}$ and $q_{1}$ and $\mathcal{F}_{{\sf CZ}(1,2)}=96.4(2)\%$ for qubit pair $q_{1}$ and $q_{2}$. The fidelity difference observed between the two pairs is attributable to position-dependent intensity variation in the Rydberg beams. Each pair of atoms are illuminated by different portions of the beam. The $q_{1}-q_{2}$ atom pair experiences faster dephasing than $q_{0}-q_{1}$ due to a less flat intensity distribution when the atoms are horizontally spaced compared to when the atoms are vertically spaced.
We used standard single-qubit Clifford randomized benchmarking [40] to characterize both global and local single-qubit gates (see Fig. 13c for a typical circuit). For global single-qubit rotations, we apply a global variant of the standard randomized benchmarking as shown in [41] (see Fig.13d). Using Eq. (20) with $d=2$, we extracted an average gate fidelity of 0.99989(1) The faster decay after 500 quantum gates is due to changes in the qubit frequency due to a transient increase of the magnetic field. A similar RB procedure was used to find the gate fidelity of a set of local Clifford gates constructed using global $\sf{R}_{\phi}$ and local $\sf{R}_{z}$ gates (see Fig.13e). We measured an average local Clifford gate fidelity of 0.99941(4).
@@ Appendix H Optimized $\sf CZ$ gate
In this section, we present the experimental modifications we made to improve the $\sf CZ$ gate fidelity on the $q_{1}$ and $q_{2}$ atom pair. These modifications were made after the VQE measurements were performed and highlight some strategies for fidelity improvement. These modifications include: 1. We excited to a lower Rydberg level, $55s_{1/2}$, which has a larger dipole matrix element compared to $66s_{1/2}$. The larger matrix element reduces the intensity required to drive the Rydberg transition at a given Rydberg Rabi rate. This effectively reduces infidelity caused by AC Stark shift fluctuations resulting from intensity fluctuations. 2. We found the optimal Rydberg Rabi frequency to be $2\pi\times 1$ MHz (versus $2\pi\times 1.7$ MHz in the VQE experiment). Although gates performed at lower Rabi frequency are more susceptible to some error sources such as Rydberg decay, the increase in such error sources is more than compensated for by the reduction in dephasing during the Rydberg excitation that was observed in the experiment. This dephasing reduction is not fully understood, but we believe that it is related to intensity noise found in the $459$-nm Rydberg laser. Additional measurements and simulations are needed to confirm this hypothesis. 3. After optical pumping, we added a trap power ramp to reduce the depth from $600\mu$K to $260~{}\mu$K. This ramping, along with improved cooling parameters, reduced the atom temperature to $6~{}\mu$K. After these changes, we measured a $\sf CZ$ gate fidelity of $\mathcal{F}_{\sf CZ}=0.986(1)$ (see Fig.14) using the methods described in Appendix G.